OpenStudy (anonymous):

Solvee The Equation Beloww ; 10 + v - 17v = 4

OpenStudy (nottim):

Meanie.

OpenStudy (nottim):

It's algebra. Rearrange and solve. Use either the BEDMAS or PEDMAS rules.

OpenStudy (anonymous):

Ok ? lol

OpenStudy (nottim):

Don't get us to do your homework. That's rude.

jhonyy9 (jhonyy9):

16v=6 v=6/16 = 3/8